OpenStudy (anonymous):

Find the value of f ′(0) for f(x) = sin(x) + e^x.

OpenStudy (er.mohd.amir):

diff. with respect to x f ' (x) = cos x+ e^x at x=0 f ' (0) = cos 0 +e^0=1+1=2

OpenStudy (anonymous):

Okay thank you!

OpenStudy (xapproachesinfinity):

don't give answer dear, read the open study code of conduct!!

OpenStudy (xapproachesinfinity):

you can only guide the poster not giving the answers